7x-8y=-2
x-3y= 9 Solving Systems of Equations by Substitution

Found 3 solutions by Boreal, stanbon, macston:
Answer by Boreal(15235)

7x-8y=-2
x-3y= 9 ; x=9 +3y
7(9+3y)-8y= -2
63+21y-8y= -2
13y= -65
y=-5
x=-6, from x=9+3y
Check by trying in first equation.
7(-6)-8(-5)=-42+40 = -2

Answer by stanbon(75887)

7x-8y=-2
x-3y= 9
------
Solve the 2nd equation for "x"::
x = 3y+9
----
Substitute for "x" and solve for "y":
7(3y+9) - 8y = -2
21y + 63 - 8y = -2
13y = -65
y = -5
------
Solve for "x"::
x = 3y + 9
x = 3*-5+9
x = -15+9
x = -6
------------
Cheers,
Stan H.
--------------

Answer by macston(5194)

.
x-3y=9
x=3y+9 . Use this to substitute for x in the other equation.
.
7x-8y=-2
7(3y+9)-8y=-2
21y+63-8y=-2
13y+63=-2
13y=-65
y=-5
.
x-3y=9
x-3(-5)=9
x-(-15)=9
x+15=9
x=-6
.
ANSWER: (-6,-5)
.
CHECK:
7x-8y=-2
7(-6)-8(-5)=-2
-42-(-40)=-2
-42+40=-2
-2=-2
.